Given the consumption of two goods X and Y subject to their respective prices and a given income graphically derive the demand curve for good X in the case of price fall.
1
Expert's answer
2020-02-12T09:09:57-0500

Solution

Let M=100, "p_y=20, Q_y=\frac {100}{20}=5"

"p_{x_1}=10, Q_{x_1}=\frac{100}{10}=10"

Then will be graph




If "p_{x_2}=5" then "Q_{x_2}=\frac {100}{5}=20"

Will be next graph